OpenStudy (anonymous):

If the endpoints of AB¯¯¯¯¯ have the coordinates A(9, 8) and B(-1, -2), what is the midpoint of AB¯¯¯¯¯? (5, 3) (4, 5) (5, 5) (4, 3)

OpenStudy (anonymous):

\[\left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\]

OpenStudy (anonymous):

that will give you your midpoint (x,y)

OpenStudy (anonymous):

is it 4,3 ? @iambatman

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

thank You @iambatman

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

A quadrilateral has vertices A(3, 5), B(2, 0), C(7, 0), and D(8, 5). Which statement about the quadrilateral is true? ABCD is a parallelogram with non-perpendicular adjacent sides. ABCD is a trapezoid with only one pair of parallel sides. ABCD is a rectangle with non-congruent adjacent sides. ABCD is a rhombus with non-perpendicular adjacent sides.